th sum of four consecutive whole numbers is 150.
a- let x represent the first number. Write expressions to represent the other three numbers
b-write and solve an equation to find the four numbers.

Answer by checkley77(12569)

lET X, X+1, X+2, X+3 BE TH 4 NUMBERSS.
X+X+1+X+2+X+3=150
4X+6=150
4X=150-6
4X=144
X=144/4
X=36 ANS. FOR THE SMALLEST NUMBER.
36+1=37 ANS. FOR THE SECOND NUMBER.
36+2=38 ANS. FOR THE THIRD NUMBER.
36=3=39 ANS. FOR THE FOURTH NUMBER.
PROOF:
36+37+38+39=150
150=150